XYZ common stock is expected to pay a dividend of $1.46 next year, and that dividend grows at a constant rate of 11.4. If the current price of XYZ common stock is $107.76, then what is the expected rate of return for this stock, based on the Discounted Cash Flow model? (Show your answer in DECIMAL FORM to three decimals, e.g., 12.3% would be entered as 0.123).
Answer #1

Current price of a stock =D1/Ke-G

D1 = dividend in year 1, that means next year

Ke = Expected rate of return

G = Growth

107.76 = 1.46/Ke - 0.114

Ke - 0.114 = 1.46/107.76

Ke = 0.0135 + 0.114

Ke = 0.1275 (or) 0.128